OpenStudy (anonymous):

How would you sketch a graph that is an even degree function, but is not an even function?

OpenStudy (anonymous):

what does that even mean? (x-5)^2 = y

OpenStudy (amistre64):

an even function is an even degree; but perhaps you are talking about rational expressions?

OpenStudy (anonymous):

does y = x^2 have an even degree?

OpenStudy (amistre64):

y = x^2 is an even degree :)

OpenStudy (anonymous):

y = x^2 + x is of degree 2, so it is even, but y is not an even function

OpenStudy (amistre64):

or do we mean even function as in "spins around the y axis"?

OpenStudy (anonymous):

why are you trying to sketch this thing? What math class are you in?

OpenStudy (anonymous):

was your last question a comment or a statement?

OpenStudy (anonymous):

an even degree implies f(-x) = f(x), or reflects across the y axis

OpenStudy (anonymous):

i'm sorry an even funciton

OpenStudy (anonymous):

so you can have even degreed functions even, or not.. ie y = x^2 vs y = (x-2)^2 which is not centered at the origin

OpenStudy (anonymous):

an even degree is y = x ^2n where n is an integer

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

(x-a)^2n involves a translation of a units to the right

OpenStudy (anonymous):

(x-a)^2n + b involves the same a translation, with a vertical translation b units (up)